# Quantum Entanglement Autodistillation in Baryon Pair Decays

### v1

@@ Abstract
We study the spin-entangled mixed state of a spin-1/2 baryon–antibaryon pair produced in the process $e^{+}e^{-}\to J/\psi,\psi^{\prime}\to\mathcal{B}\bar{\mathcal{B}}$. We show that the spin entanglement of the system can increase following the decays $\mathcal{B}\to\mathcal{b}+M$ and ¯$\bar{\mathcal{B}}\to\bar{\mathcal{b}}+\bar{M}$, where $\mathcal{b},\bar{\mathcal{b}}$ are spin-1/2 baryons and $M,\bar{M}$ are spin-0 mesons. This phenomenon, known as entanglement autodistillation, manifests as a probabilistic amplification of entanglement during the decay process. We analyze the underlying mechanism and show that this phenomenon depends only on the initial state of the $\mathcal{B}\bar{\mathcal{B}}$ system and the decay parameter $\alpha_{D}$, but not on the decay parameter $\phi_{D}$.
@@ 1 Introduction
The application of quantum entanglement concepts to particle physics has become a growing area of interest in recent years[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. A recent study[24] reported the observation of quantum entanglement in top-quark pairs at the LHC using the ATLAS detector. This represents the highest energy scale at which entanglement has been observed to date. The top quark, the heaviest known elementary particle, has an extremely short lifetime ($10^{-25}s$), significantly shorter than the hadronization timescale. Theoretical studies on top-quark entanglement have made substantial progress [25, 26, 27, 28], and experimental observations hold significant implications: on one hand, they provide a test of fundamental quantum entanglement concepts at high energy and ultrashort timescales; on the other hand, entanglement may offer novel methods for probing particle physics. This research opens new perspectives and opportunities at the intersection of quantum information and quantum field theory.
Recent studies [29, 30, 31] have shown that particle decay processes can significantly change a system’s entanglement, potentially leading to an anomalous phenomenon known as “entanglement amplification”. This process, referred to as “autodistillation”, suggests that entanglement can spontaneously increase during the system’s evolution, although with some probability of failure. According to quantum information principles, entanglement cannot increase through local operations and classical communication (LOCC). However, under suitable local manipulations, it may increase probabilistically. These operations, known as stochastic LOCC (SLOCC) [32], form the theoretical basis of autodistillation. The transformation of entanglement in two-qubit mixed states under SLOCC operations has been studied in Ref. [33], where the so-called “filtering operations” are identified as SLOCC transformations.
@@ 3 Specific Process
  • ¯$\bar{\mathcal{B}}\to\bar{\mathcal{b}}+\bar{M}_{0}$ (antibaryon decay)
@@
The complete process can be characterized by kinematic variables $\xi=(\theta_{1},\theta_{\mathcal{b}},\phi_{\mathcal{b}},\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})$, where $\theta_{1}$ represents represents the scattering angle of the $\mathcal{B}$ baryon with respect to the positron beam in the center-of-mass (CM) frame. The remaining four angles, $\theta_{\mathcal{b}},\phi_{\mathcal{b}},\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}}$, describe the orientation of the $\mathcal{b}$ and $\bar{\mathcal{b}}$ baryons in their respective helicity frames, denoted as $\mathcal{R}_{\mathcal{B}}$ and $\mathcal{R}_{\bar{\mathcal{B}}}$, where the $\mathcal{B}$ and $\bar{\mathcal{B}}$ baryons are at rest.
@@
in which the $\sigma^{\mathcal{B}}_{\mu}$ matrices are in the mother helicity frame and $\sigma_{\nu}^{\mathcal{b}}$ in the daughter helicity frame. The elements of the matrix $a_{\mu\nu}$ are parameterized by the decay parameters $\alpha_{D}$ and $\phi_{D}$ and depend on the decay angles, i.e. $a_{\mu\nu}^{\mathcal{B}}(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\alpha_{D}^{\mathcal{B}},\phi_{D}^{\mathcal{B}})$. The decay parameters $\alpha_{D}$ and $\phi_{D}$ are defined in [40]. The decay parameter $\phi_{D}$, constrained within $[-\pi,\pi]$, is related to the rotation of the spin vector between mother baryon $\mathcal{B}$ and daughter baryon $\mathcal{b}$. The decay parameter $\alpha_{D}$, constrained within $[-1,1]$, is determined from the angular distribution asymmetry of the $\mathcal{B}$ baryon in the $\mathcal{B}$ baryon rest frame [41]. Therefore, the density matrix of daughter baryon and antibaryon would read:
@@ 4 Results
Following the methodology outlined in Section 3, we now present the comparison between $\mathcal{N}(\rho_{\mathcal{B}\bar{\mathcal{B}}})$ and $\mathcal{N}(\rho_{\mathcal{b}\bar{\mathcal{b}}})$(and similarly for $\mathscr{C}$). The decay angle $(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})$ of baryon and antibaryon in their helicity frame have been adjusted to find the most appropriate final density matrix for observing the phenomenon where entanglement increases after decay.
@@ 4.1 Analytic result of a hypothetical example
We will show how the state after decay can be more entangled than before with a special hypothetical example. For simplicity, the real parameters $\alpha_{\psi}$ and $\Delta\Phi$ describing the initial baryon-antibaryon state are assumed to be 00 and $\pi/2$, the corresponding initial density matrix is
@@
The decay parameters $\alpha_{D}^{\mathcal{B}}$ and $\phi_{D}^{\mathcal{B}}$ of baryon decay process are assumed to be $1/2$ and $\pi/2$. We assume that there is no charge-parity symmetry breaking in this process. Therefore, the antibaryon decay parameters $\alpha_{D}^{\bar{\mathcal{B}}}$ and $\phi_{D}^{\bar{\mathcal{B}}}$ are assumed to be $-1/2$ and $\pi/2$. We choose $\phi_{\mathcal{b}}=\phi_{\bar{\mathcal{b}}}=\pi/2$, which denotes that the plane of the final baryon-antibaryon pair is perpendicular to the scattering plane of $e^{+}e^{-}\to\mathcal{B}\bar{\mathcal{B}}$ process. Consider the situation where the scattering angle is $\theta_{1}=\pi/4$, and the concurrence of the initial state is $1/4$. In this situation, the full expression of the concurrence of the final baryon-antibaryon pair $\mathcal{b}\bar{\mathcal{b}}$, which, depending on the decay angles $\theta_{\mathcal{b}}$ and $\theta_{\bar{\mathcal{b}}}$, is
@@ 4.2 Baryon and antibaryon decay
Figure 2: (a): Negativity of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{-}+\bar{\Xi}^{+}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b): Concurrence of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{-}+\bar{\Xi}^{+}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
As shown in Fig. 2, negativity and concurrence depend on the scattering angle $\theta_{1}$ of the $e^{+}e^{-}\to\Xi^{-}\bar{\Xi}^{+}$ process. Our calculation of the concurrence of the $\Xi^{-}\bar{\Xi}^{+}$ state is consistent with the result reported in Ref. [6]. We note that the helicity state of $\Xi^{-}\bar{\Xi}^{+}$ is a separable mixed state (i.e. $\rho=\sum_{i}p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)}$ and there is no entanglement) when $\theta_{1}=0\text{ or }\pi$. The final state of $\Lambda\bar{\Lambda}$ is also a separable mixed state if there is no entanglement of $\Xi^{-}\bar{\Xi}^{+}$. When $\theta_{1}=\pi/2$, although the concurrence and negativity of $\rho_{\Xi^{-}\bar{\Xi}^{+}}$ almost reach the maximum, they are not increase at this point. This phenomenon is much more apparent in Fig. 3 as shown below.
@@
Figure 3: (a): increase of negativity after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b): increase of concurrence after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 4: (a): Negativity of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to\psi(3686)\to\Xi^{-}+\bar{\Xi}^{+}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b): Concurrence of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to\psi(3686)\to\Xi^{-}+\bar{\Xi}^{+}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 5: (a): increase of negativity after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b): increase of concurrence after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 6: (a): Negativity of $\Xi^{0}\bar{\Xi}^{0}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{0}+\bar{\Xi}^{0}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize negativity. (b): Concurrence of $\Xi^{0}\bar{\Xi}^{0}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{0}+\bar{\Xi}^{0}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 7: (a): increase of negativity after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize negativity. (b): increase of concurrence after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize concurrence.

### v2

@@ Abstract
We study the spin-entangled mixed state of a spin-1/2 baryon–antibaryon pair produced in the process $e^{+}e^{-}\to J/\psi,\psi^{\prime}\to\mathcal{B}\bar{\mathcal{B}}$. We show that the spin entanglement of the system can increase following the decays $\mathcal{B}\to\mathcal{b}+M$ and $\bar{\mathcal{B}}\to\bar{\mathcal{b}}+\bar{M}$, where $\mathcal{b},\bar{\mathcal{b}}$ are spin-1/2 baryons and $M,\bar{M}$ are spin-0 mesons. This phenomenon, known as entanglement autodistillation, manifests as a probabilistic amplification of entanglement during the decay process. We analyze the underlying mechanism and show that this phenomenon depends only on the initial state of the $\mathcal{B}\bar{\mathcal{B}}$ system and the decay parameter $\alpha_{D}$, but not on the decay parameter $\phi_{D}$. This effect only arises when parity is violated in the decay, when $\alpha_{D}\neq 0$.
@@ 1 Introduction
The application of quantum entanglement concepts to particle physics has become a growing area of interest in recent years[2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. A recent study[27] reported the observation of quantum entanglement in top-quark pairs at the LHC using the ATLAS detector. This represents the highest energy scale at which entanglement has been observed to date. The top quark, the heaviest known elementary particle, has an extremely short lifetime ($10^{-25}s$), significantly shorter than the hadronization timescale. Theoretical studies on top-quark entanglement have made substantial progress [28, 29, 30, 31], and experimental observations hold significant implications: on one hand, they provide a test of fundamental quantum entanglement concepts at high energy and ultrashort timescales; on the other hand, entanglement may offer novel methods for probing particle physics. This research opens new perspectives and opportunities at the intersection of quantum information and quantum field theory.
Recent studies [32, 33, 34, 35] have shown that particle decay processes can significantly change a system’s entanglement, potentially leading to an anomalous phenomenon known as “entanglement amplification”. This process, referred to as “autodistillation”, suggests that entanglement can spontaneously increase during the system’s evolution, although with some probability of failure. According to quantum information principles, entanglement cannot increase through local operations and classical communication (LOCC). However, under suitable local manipulations, it may increase probabilistically. These operations, known as stochastic LOCC (SLOCC) [36], form the theoretical basis of autodistillation. The transformation of entanglement in two-qubit mixed states under SLOCC operations has been studied in Ref. [37], where the so-called “filtering operations” are identified as SLOCC transformations.
@@ 3 Specific Process
  • $\bar{\mathcal{B}}\to\bar{\mathcal{b}}+\bar{M}_{0}$ (antibaryon decay)
@@
The complete process can be characterized by kinematic variables $\xi=(\theta_{1},\theta_{\mathcal{b}},\phi_{\mathcal{b}},\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})$, where $\theta_{1}$ represents the scattering angle of the $\mathcal{B}$ baryon with respect to the positron beam in the center-of-mass (CM) frame. The remaining four angles, $\theta_{\mathcal{b}},\phi_{\mathcal{b}},\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}}$, describe the orientation of the $\mathcal{b}$ and $\bar{\mathcal{b}}$ baryons in their respective helicity frames, denoted as $\mathcal{R}_{\mathcal{B}}$ and $\mathcal{R}_{\bar{\mathcal{B}}}$, where the $\mathcal{B}$ and $\bar{\mathcal{B}}$ baryons are at rest.
@@
in which the $\sigma^{\mathcal{B}}_{\mu}$ matrices are in the mother helicity frame and $\sigma_{\nu}^{\mathcal{b}}$ in the daughter helicity frame. The elements of the matrix $a_{\mu\nu}$ are parameterized by the decay parameters $\alpha_{D}$ and $\phi_{D}$ and depend on the decay angles, i.e. $a_{\mu\nu}^{\mathcal{B}}(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\alpha_{D}^{\mathcal{B}},\phi_{D}^{\mathcal{B}})$. The decay parameters $\alpha_{D}$ and $\phi_{D}$ are defined in [44]. The decay parameter $\phi_{D}$, constrained within $\pi,\pi]$, is related to the rotation of the spin vector between mother baryon $\mathcal{B}$ and daughter baryon $\mathcal{b}$. The decay parameter $\alpha_{D}$, constrained within [1,1][-1,1], is determined from the angular distribution asymmetry of the $\mathcal{B}$ baryon in the $\mathcal{B}$ baryon rest frame [45]. Therefore, the density matrix of daughter baryon and antibaryon would read:
@@ 4 Results
Following the methodology outlined in Section 3, we now present the comparison between $\mathcal{N}(\rho_{\mathcal{B}\bar{\mathcal{B}}})$ and $\mathcal{N}(\rho_{\mathcal{b}\bar{\mathcal{b}}})$(and similarly for $\mathscr{C}$). We provide an analytical result from a hypothetical example to demonstrate entanglement autodistillation in Section 4.1. In Section 4.2, we numerically simulate several specific physical processes to observe this phenomenon. In these simulations, the decay angles $(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})$ of baryon and antibaryon in their helicity frame have been adjusted to find the most appropriate final density matrix for observing the phenomenon where entanglement increases after decay.
@@ 4.1 Analytic result of a hypothetical example
We will show how the state after decay can be more entangled than before with a special hypothetical example. For simplicity, the real parameters $\alpha_{\psi}$ and $\Delta\Phi$ describing the initial baryon-antibaryon state are assumed to be 0 and $\pi/2$, the corresponding initial density matrix is
@@
The decay parameters $\alpha_{D}^{\mathcal{B}}$ and $\phi_{D}^{\mathcal{B}}$ of baryon decay process are assumed to be $1/2$ and $\pi/2$. We assume that there is no charge-parity symmetry breaking in this process. Therefore, the antibaryon decay parameters $\alpha_{D}^{\bar{\mathcal{B}}}$ and $\phi_{D}^{\bar{\mathcal{B}}}$ are assumed to be 1/2-1/2 and $\pi/2$. We choose $\phi_{\mathcal{b}}=\phi_{\bar{\mathcal{b}}}=\pi/2$, which denotes that the plane of the final baryon-antibaryon pair is perpendicular to the scattering plane of $e^{+}e^{-}\to\mathcal{B}\bar{\mathcal{B}}$ process. Consider the situation where the scattering angle is $\theta_{1}=\pi/4$, and the concurrence of the initial state is $1/4$. In this situation, the full expression of the concurrence of the final baryon-antibaryon pair $\mathcal{b}\bar{\mathcal{b}}$, which, depending on the decay angles $\theta_{\mathcal{b}}$ and $\theta_{\bar{\mathcal{b}}}$, is
@@ 4.2 Baryon and antibaryon decay
Figure 2: (a) Negativity of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{-}+\bar{\Xi}^{+}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b) Concurrence of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{-}+\bar{\Xi}^{+}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
As shown in Fig. 2, negativity and concurrence depend on the scattering angle $\theta_{1}$ of the $e^{+}e^{-}\to\Xi^{-}\bar{\Xi}^{+}$ process. Our calculation of the concurrence of the $\Xi^{-}\bar{\Xi}^{+}$ state is consistent with the result reported in Ref. [6]. We note that the helicity state of $\Xi^{-}\bar{\Xi}^{+}$ is a separable mixed state (i.e. $\rho=\sum_{i}p_{i}\rho_{i}^{(1)}\otimes\rho_{i}^{(2)}$ and there is no entanglement) when $\theta_{1}=0\text{ or }\pi$. The final state of $\Lambda\bar{\Lambda}$ is also a separable mixed state if there is no entanglement of $\Xi^{-}\bar{\Xi}^{+}$. When $\theta_{1}=\pi/2$, although the concurrence and negativity of $\rho_{\Xi^{-}\bar{\Xi}^{+}}$ almost reach the maximum, they do not increase at this point. This phenomenon is much more apparent in Fig. 3 as shown below.
@@
Figure 3: (a) Increase of negativity after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b) Increase of concurrence after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 4: (a) Negativity of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to\psi(3686)\to\Xi^{-}+\bar{\Xi}^{+}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b) Concurrence of $\Xi^{-}\bar{\Xi}^{+}$ from $e^{+}e^{-}\to\psi(3686)\to\Xi^{-}+\bar{\Xi}^{+}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
As shown in Fig. 4, the entanglement behavior of the process $\psi(3686)\to\Xi^{-}+\bar{\Xi}^{+}$, followed by the decays $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$, exhibits a pattern similar to that seen in Fig. 2. The negativity and concurrence of the $\Xi^{-}\bar{\Xi}^{+}$ state depend on the scattering angle $\theta_{1}$, and the entanglement vanishes at $\theta_{1}=0$ or $\pi$, where the helicity state becomes separable. Our results for the initial-state concurrence are consistent with those reported in Ref. [6]. Compared to the $J/\psi$ channel in Fig. 2, the increase of entanglement after decay is less pronounced, as can be seen in Fig. 5. Additionally, Fig. 5 exhibits larger uncertainties in the post-decay entanglement, mainly due to the uncertainties in the input parameters $\alpha$ and $\Delta\Phi$.
@@
Figure 5: (a) Increase of negativity after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize negativity. (b) Increase of concurrence after $\Xi^{-}\to\Lambda\pi^{-}$ and $\bar{\Xi}^{+}\to\bar{\Lambda}\pi^{+}$ with the most appropriate decay angle to maximize concurrence.
@@
Figure 6: (a) Negativity of $\Xi^{0}\bar{\Xi}^{0}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{0}+\bar{\Xi}^{0}$ process and Negativity of $\Lambda\bar{\Lambda}$ after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize negativity. (b) Concurrence of $\Xi^{0}\bar{\Xi}^{0}$ from $e^{+}e^{-}\to J/\psi\to\Xi^{0}+\bar{\Xi}^{0}$ process and Concurrence of $\Lambda\bar{\Lambda}$ after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize concurrence.
As shown in Fig. 6, both the negativity and concurrence of the $\Xi^{0}\bar{\Xi}^{0}$ state produced in the process $e^{+}e^{-}\to J/\psi\to\Xi^{0}+\bar{\Xi}^{0}$ also exhibit dependence on the scattering angle $\theta_{1}$. Similar to the charged case, our results are consistent with those reported in Ref. [6]. In particular, the entanglement vanishes when $\theta_{1}=0$ or $\pi$, corresponding to a separable helicity state. Consequently, the final state of $\Lambda\bar{\Lambda}$, produced via $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$, remains unentangled under these conditions. At $\theta_{1}=\pi/2$, the entanglement of the $\Xi^{0}\bar{\Xi}^{0}$ state nearly reaches its maximum; however, the entanglement of the final state does not increase at this point, a feature that is clearly illustrated in Fig. 7.
@@
Figure 7: (a) Increase of negativity after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize negativity. (b) Increase of concurrence after $\Xi^{0}\to\Lambda\pi^{0}$ and $\bar{\Xi}^{0}\to\bar{\Lambda}\pi^{0}$ with the most appropriate decay angle to maximize concurrence.
As discussed earlier, the decay angles $(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})$ of the baryon and antibaryon in their respective helicity frames have been adjusted to identify the optimal final density matrix that reveals the entanglement amplification phenomenon. In most cases, multiple sets of decay angles can lead to the maximum value of entanglement after the decay process. However, in the scenarios shown in Sections 4.2.1 and 4.2.3, there exist unique decay angles corresponding to specific scattering angle intervals, as revealed by numerical simulations. Specifically, when the scattering angle $\theta_{1}\in[0.22\pi,0.42\pi]$, the unique configuration is
$(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})=\left(\tfrac{1}{2}\pi,\tfrac{3}{2}\pi;\tfrac{1}{2}\pi,\tfrac{3}{2}\pi\right),$
and when $\theta_{1}\in[0.58\pi,0.78\pi]$, the unique configuration becomes
$(\theta_{\mathcal{b}},\phi_{\mathcal{b}};\theta_{\bar{\mathcal{b}}},\phi_{\bar{\mathcal{b}}})=\left(\tfrac{1}{2}\pi,\tfrac{1}{2}\pi;\tfrac{1}{2}\pi,\tfrac{1}{2}\pi\right).$
Fortunately, as illustrated in Fig. 3 and Fig. 7, these regions coincide with the optimal conditions for observing entanglement enhancement after decay.